# Classification of unimodal isolated complete intersection singularities in positive characteristic

### v2

@@ 2.5. Modality
V. I. Arnol’d introduced the notion of modality in his famous [AVGZ12] as follows: The modality of a point $x\in X$ under the action of a Lie group $G$ on a manifold $X$ is the smallest $m$ such that a sufficiently small neighborhood of $x$ may be covered by a finite number of orbit families of $m$ parameters. G.-M. Greuel and Nguyen H. D. generalize the notion and give a detailed discuss in [GN16], [Ngu13]. For the definition of the modality of an ICIS, we refer to [PPG24] Remark 1.13(3).
@@
First we give a lower bound by complete transversal in [BAM20], which is useful in next section.
@@
(2.5) $\mathrm{cod}(f+a)=\mathrm{comdimension\ of\ }\widetilde{T}_{f}(\mathcal{K}_{l}f)\cap P_{k,l}\mathrm{\ in\ }P_{k,l}$
@@
(2.6) $\mathrm{cod}_{0}(f)=\mathrm{inf}_{a\in C}\{\mathrm{cod}(f+a)\}.$
@@
Let $f\in J_{k}$ be a $k$-jet of weighted homogeneous type with degree 0 w.r.t. $(a_{1},\dots,a_{n};d_{1},\dots,d_{m})$. Then for $a\in U,$ $f+a$ has modality $\mathrm{cod}_{0}(f)$ in $J_{l}(f)$ under the action of the subgroup $\mathcal{K}_{l}(f)$ of $\mathcal{K}_{l}$ which stabilize $f$. In particular, any jet in $J_{l}(f)$ has $\mathcal{K}_{l}(f)$modality $\geq\mathrm{cod}_{0}(f)$ in $J_{l}$.
@@
The idea mainly comes from [BAM20] Proposition 1.4. We rewrite the proof using tangent image instead of tangent space for the sake of fields with positive characteristic.
Find a subspace $\langle e_{1},\dots,e_{c}\rangle\in P_{k,l}$ with $\langle e_{1},\dots,e_{c}\rangle\oplus\widetilde{T}_{f}(\mathcal{K}_{l}(f+a))\cap P_{k,l}=P_{k,l}$. Then $\mathrm{cod}(f+a)=c.$ Since $\widetilde{T}_{f}(\mathcal{K}_{l}(f+b))$ varies continuously for $b\in U$, we have $\langle e_{1},\dots,e_{c}\rangle\cap\widetilde{T}_{f}(\mathcal{K}_{l}(f+b))=\{0\}$ for $b$ in an Zariski open neighborhood $V$ of $a$.
@@
(2.7) $\phi:(F^{c},0)\rightarrow(J_{l}(f),f+a),\ (t_{1},\dots,t_{c})\mapsto f+a+\sum_{i=1}^{c}t_{i}e_{i}.$
@@
For any $g\in J_{l}(f),$ write $g=f+\widetilde{g}=(f_{1}+\widetilde{g_{1}},\dots,f_{m}+\widetilde{g_{m}}),$ where $\widetilde{g}=(\widetilde{g_{1}},\dots,\widetilde{g_{m}})$ with weighted degree of $\widetilde{g_{i}}>d_{i}.$ We have
(2.8) $g(x)\sim g_{t}(x)$ $=(t^{-d_{1}}g_{1}(t_{1}^{a_{1}}x_{1},\dots,t_{n}^{a_{n}}x_{n}),\dots,t^{-d_{n}}g_{m}(t_{1}^{a_{1}}x_{1},\dots,t_{n}^{a_{n}}x_{n}))$
$=(f_{1}+t^{-d_{1}}\widetilde{g_{1}}(t_{1}^{a_{1}}x_{1},\dots,t_{n}^{a_{n}}x_{n}),\dots,f_{m}+t^{-d_{n}}\widetilde{g_{m}}(t_{1}^{a_{1}}x_{1},\dots,t_{n}^{a_{n}}x_{n}))$
$g(x)\sim g_{t}(x)=(t^{-d_{1}}g_{1}(t_{1}^{a_{1}}x_{1},\dots,t_{n}^{a_{n}}x_{n}),\dots,t^{-d_{n}}g_{m}(t_{1}^{a_{1}}x_{1},\dots,t_{n}^{a_{n}}x_{n})).$
Since $f$ is weighted homogeneous w.r.t. $(a;d)$, we may furthermore choose suitable $a,d$ such that
(2.9) $\mathrm{max}(d_{i})<(k+1)\mathrm{min}(a_{j})\mathrm{\ or\ }\mathrm{min}(d_{i})>(k+1)\mathrm{max}(a_{j})$
Then condition 2.9 ensures any neighborhood of $f$ intersects $\mathcal{K}_{l}(f)\cdot g$, that is, any neighborhood of $f$ intersects all $\mathcal{K}_{l}(f)$-orbits in $J_{l}(f).$ Hence for any $\mathcal{K}_{l}(f)$-orbit $X$ in $J_{l}(f)$, there exists $b\in V,\ X=\mathcal{K}_{l}(f)\cdot(f+b).$ But we have $\widetilde{T}_{f+b}(\mathcal{K}_{l}(f)\cdot(f+b))=\widetilde{T}_{f+b}(\mathcal{K}_{l}\cdot(f+b))\cap P_{n,l},$ hence $\widetilde{T}_{f+b}(\mathcal{K}_{l}(f)\cdot(f+b))\cap\langle e_{1},\dots,e_{c}\rangle=\{0\}$, i.e. $\phi^{-1}(X)$ has only finitely many points in a neighborhood of 0. This finishes the proof. ∎
@@
while for any $g\in X^{\prime}$, $\mathrm{dim}(\mathcal{K}^{\prime}\cdot g)\leq\mathrm{dim}\mathcal{K}^{\prime}-1$, since $\{(a^{l}E_{n},\frac{1}{a}E_{n})\mid a\in F^{\times}\}\subset\mathcal{K}^{\prime}$ stabilize $g$.
@@
Let $f$ be an unimodal zero-dimensional ICIS, then $f\in I_{2,2}$ or $I_{3,3}$.
@@ 3. The classification of order 2
Some classification of order 2 ICIS has already been discussed in [PPG24], i.e. ICIS of modality 0. Here we continue their work and finish the classification of order 2 ICIS with modality 1. In this section $f=(f_{1},f_{2})\in R^{2}.$
First, we assume $charF\neq 2$. The following two propositions are from [PPG24].
@@
(v) Let $t>q$ and $\alpha=1$. Then $f\sim(x^{2}+y^{s},y^{t}+exy^{q})$ for a suitable unit $e\in F[[y]]$. If $\ 2t-2q-s\neq 0$ and ( $p=0$ or $p\nmid 2t-2q-s$ ), then $f\sim(x^{2}+y^{s},y^{t}+xy^{q})$.
@@
Assume $f$ is of the form 3.1. Then if $s\geq 5,t\geq 6$ and $q\geq 3$, then $f$ is of modality at least 2.
@@
Let $g=j_{3}(f)=(x^{2},0).$ Then any 6-jet $h\in J_{6}$ with $j_{4}(h)=g$ are equivalent to $h\sim(x^{2}+\alpha y^{5},y^{6}+axy^{3}+bxy^{4})$. For all $a,b\in F$, computation shows $(y^{4},0),(0,y^{4}),(0,y^{5})\notin\widetilde{T}_{h}(\mathcal{K}_{6}h)$. Hence the codimension of $\widetilde{T}_{h}(\mathcal{K}_{6}h)\cap P_{4,6}$ in $P_{4,6}$ is at least 2. By Proposition 2.19, $\mathcal{K}-mod(f)\geq\mathcal{K}_{6}-mod(f)\geq 2$. ∎
@@
By 3.1, $h\sim(x^{2}+\alpha y^{5},y^{6}+axy^{3}+bxy^{4}+cxy^{5})$. One can show, for example, applying $\phi(y)=y-\frac{c}{6}x$, that $h\sim(x^{2}+\alpha y^{5},y^{6}+axy^{3}+bxy^{4})$. This simplifies the computation of the codimension, but does not affect the result.
The following proposition is also from [PPG24].
@@
The following ICIS are the only candidates for being simple (i.e of modality 0):
@@
            (3.b.ii.2.1) $s\geq 3,\ q\leq 2,$ then $f\sim(xy,x^{2}+y^{m})$ for some $m$ for $q=1$, and $f\sim(x^{2}+y^{m},xy^{2})$ for some $m$ for $q=2$.
@@
$\widetilde{l}_{q,t}$ $(x^{2}+y^{4},y^{q+2}+\lambda xy^{q}+uxy^{t}+xy^{t^{\prime}}),\ where$ $\lambda^{2}=-1,q\geq 3,$
@@
               (3.b.i.2.2.3) If $p=t=3,\ s>4,$ same as the process in [PPG24] Proposition 2.5, $f\sim(x^{2},y^{3}+xy^{2})$, which is simple.
               (3.b.i.2.2.4) If $p=t=5,\ s>4$, then
(3.2) $f$ $\sim(x^{2}+y^{s},y^{5}+exy^{4})$
$\sim(x^{2}+e_{0}y^{s},y^{5}+xy^{4})$
@@
where $\widetilde{e}=\frac{1}{1+\frac{e_{0}}{2}y^{s-2}}$. Applying $\phi(x)=\frac{1}{\widetilde{e}}x$ and $\phi(y)=y$ , we have $f\sim(x^{2}+e_{1}y^{2s-1},y^{5}+xy^{4}).$ Repeating the process, we get $f\sim(x^{2},y^{5}+xy^{4})$ with $p=charF=5$.
@@
         (3.b.ii.1) If $\alpha=0$, by Proposition 3.3, we have $1\geq q\geq 2$ or $q=3,t=5$.
@@
            (3.b.ii.2.2) If $s=3$, then $f\sim(x^{2}+y^{3},xy^{q})$ or $(x^{2}+y^{3},y^{q+2}+xy^{q})$, which is also simple.
@@
               (3.b.ii.2.3.1) If $t\geq q+4$, then $f\sim(x^{2}+y^{4},y^{t}+exy^{1}-y^{t-4}(x^{2}+y^{4}))=(x^{2}+y^{4},(e-xy^{t-q-4})xy^{q})\sim(x^{2}+y^{4},xy^{q}).$
               (3.b.ii.2.3.2) If $t=q+3$, then $f\sim(x^{2}+y^{4},y^{q+3}+xy^{q})$ by Proposition 3.2 (v) since in this case $22t-2q-s=$ 2.
               (3.b.ii.2.3.3) If $t=q+2$. Let $g=(x^{2}+y^{4},y^{q+2}+\lambda xy^{q})$, where $\lambda=e(0)\in F^{\times}.$ Then $g$ is weighted homogeneous of degree 0 w.r.t. $(a;d)$, where $a=(2,1),d=(4,q+2)$, and $g$ is the weighted 0-jet of $f$ w.r.t. $(a;d)$.
@@
for $\lambda^{2}+1=0,r>q$ and $e(y)$ is a unit in $F[[y]]$. In fact, we can show that for some $e(y)$ and $l$, $f\sim(x^{2}+y^{4},y^{q+2}+\lambda xy^{q}+xy^{l})$, while for others, we get a family of unimodal ICIS with the form
(3.4) $(x^{2}+y^{4},y^{q+2}+\lambda xy^{q}+u(y)xy^{t}+xy^{t^{\prime}}),$
@@
Denote $e_{1}=(x^{2},y^{4},0),e_{2}=(0,x^{2}+y^{4}),e_{3}=(y^{q+2}+\lambda xy^{q},0),e_{4}=(0,y^{q+2}+\lambda xy^{q}),e_{5}=(2x,\lambda y^{q}),e_{6}=(4y^{3},(q+2)y^{q+1}+\lambda qxy^{q-1})$.
Set the weight $weight(x)=2,weight(y)=1,weight(x^{2}+y^{4})=4,weight(y^{q+2}+\lambda xy^{q})=q+2.$ Then by 2.3, $\widetilde{T}_{g}(F^{1}\mathcal{K}g)$ has the element $x^{i}y^{j}e_{k}$ with:
(a) $k=1,2,3,4,\ weight(x^{i}y^{j}e_{k})\geq(5,q+3)$
@@
$Proof\ of\ the\ claim\ in$(3.b.ii.2.3.3.2):
We have $f\sim(x^{2}+y^{4},y^{q+2}+\lambda xy^{q}+e(y)xy^{t})$. If $p\nmid t-q$, we can use an ‘$\alpha,\beta$-trick’ based on the Implicit Function Theorem to show that $f\sim(x^{2}+y^{4},y^{q+2}+\lambda xy^{q}+xy^{t})$ for $t\geq q+1$. See Remark 5.11. For the case $p\mid t-q$, write $f\sim(x^{2}+y^{4},y^{q+2}+\lambda xy^{q}(1+e_{1}y^{p}+e_{2}y^{2p}+\dots)+xy^{t^{\prime}}(e_{0}^{\prime}+e_{1}^{\prime}y+e_{2}^{\prime}y^{2}+\dots))$ with $p\nmid t^{\prime}-q$. Therefore we can use $\alpha,\beta$-trick again to reduce $e_{0}^{\prime}+e_{1}^{\prime}y+\dots$. Then we have $f\sim(x^{2}+y^{4},y^{q+2}+\lambda xy^{q}+u(y)xy^{t}+xy^{t^{\prime}})$, where $u(y)=u_{0}+u_{1}y^{p}+u_{2}y^{2p}+\dots$ as we want.

### v3

@@ 2.5. Modality
V.I. Arnol’d introduced the notion of modality in his famous [AVGZ12] as follows: The modality of a point $x\in X$ under the action of a Lie group $G$ on a manifold $X$ is the smallest $m$ such that a sufficiently small neighborhood of $x$ may be covered by a finite number of orbit families of $m$ parameters. G.M. Greuel and H.D. Nguyen generalize the notion and give a detailed discuss in [GN16], [Ngu13]. For the definition of the modality of an ICIS, we refer to [PPG25] Remark 1.13(3).
@@
First we give a lower bound by complete transversal in [DG83], which is useful in next section.
@@
$\mathrm{cod}(f+a)=\mathrm{comdimension\ of\ }\widetilde{T}_{f}(\mathcal{K}_{l}f)\cap P_{k,l}\mathrm{\ in\ }P_{k,l}$ (2.6)
@@
$\mathrm{cod}_{0}(f)=\mathrm{inf}_{a\in C}\{\mathrm{cod}(f+a)\}.$ (2.7)
@@
Let $f\in J_{k}$ be a $k$-jet of weighted homogeneous type with degree 0 with respect to $(a_{1},\dots,a_{n};d_{1},\dots,d_{m})$ and satisfies condition (2.2). Then for $a\in U,$ $f+a$ has modality $\mathrm{cod}_{0}(f)$ in $J_{l}(f)$ under the action of the subgroup $\mathcal{K}_{l}(f)$ of $\mathcal{K}_{l}$ which stabilize $f$. In particular, any jet $h$ in $J_{l}(f)$ has $\mathcal{K}_{l}(f)-mod(h)\geq\mathrm{cod}_{0}(f)$ in $J_{l}$.
@@
The idea mainly comes from [DG83] Proposition 1.4. We rewrite the proof using tangent image instead of tangent space for the sake of fields with positive characteristic.
Find a subspace $\langle e_{1},\dots,e_{c}\rangle\in P_{k,l}$ with $\langle e_{1},\dots,e_{c}\rangle\oplus\widetilde{T}_{f}(\mathcal{K}_{l}(f+a))\cap P_{k,l}=P_{k,l}$. Then $\mathrm{cod}(f+a)=c.$ Since $\widetilde{T}_{f}(\mathcal{K}_{l}(f+b))$ varies continuously for $b\in U$, we have $\langle e_{1},\dots,e_{c}\rangle\cap\widetilde{T}_{f}(\mathcal{K}_{l}(f+b))=\{0\}$ for $b$ in a Zariski open neighborhood $V$ of $a$.
@@
$\phi:(F^{c},0)\rightarrow(J_{l}(f),f+a),\ (t_{1},\dots,t_{c})\mapsto f+a+\sum_{i=1}^{c}t_{i}e_{i}.$ (2.8)
@@
[…]
$g(x)\sim g_{t}(x)$ $=(t^{-d_{1}}g_{1}(t_{1}^{a_{1}}x_{1},\dots,t_{n}^{a_{n}}x_{n}),\dots,t^{-d_{n}}g_{m}(t_{1}^{a_{1}}x_{1},\dots,t_{n}^{a_{n}}x_{n}))$ (2.9)
${=(f_{1}+t^{-d_{1}}\widetilde{g_{1}}(t_{1}^{a_{1}}x_{1},\dots,t_{n}^{a_{n}}x_{n}),\dots,f_{m}+t^{-d_{n}}\widetilde{g_{m}}(t_{1}^{a_{1}}x_{1},\dots,t_{n}^{a_{n}}x_{n})).}$
Then condition (2.2) ensures any neighborhood of $f$ intersects $\mathcal{K}_{l}(f)\cdot g$, that is, any neighborhood of $f$ intersects all $\mathcal{K}_{l}(f)$-orbits in $J_{l}(f).$ Hence for any $\mathcal{K}_{l}(f)$-orbit $X$ in $J_{l}(f)$, there exists $b\in V,\ X=\mathcal{K}_{l}(f)\cdot(f+b).$ But we have $\widetilde{T}_{f+b}(\mathcal{K}_{l}(f)\cdot(f+b))=\widetilde{T}_{f+b}(\mathcal{K}_{l}\cdot(f+b))\cap P_{k,l},$ hence $\widetilde{T}_{f+b}(\mathcal{K}_{l}(f)\cdot(f+b))\cap\langle e_{1},\dots,e_{c}\rangle=\{0\}$, i.e. $\phi^{-1}(X)$ has only finitely many points in a neighborhood of 0. This finishes the proof. ∎
@@
while for any $g\in X^{\prime}$, $\mathrm{dim}(\mathcal{K}^{\prime}\cdot g)\leq\mathrm{dim}\mathcal{K}^{\prime}-1$, since $\{(a^{l}E_{n},\frac{1}{a}E_{n})\mid a\in F^{\times}\}\subset\mathcal{K}^{\prime}$ stabilize $g$, where $F\{$$^{\times}=F\backslash\{0\}$ denotes the units in $F$.
@@
Let $f$ be a unimodal zero-dimensional ICIS, then $f\in I_{2,2}$ or $I_{3,3}$.
@@ 3. The classification of order 2
Some classification of order 2 ICIS has already been discussed in [PPG25], i.e. ICIS of modality 0. Here we continue their work and finish the classification of order 2 ICIS with modality 1. In this section $f=(f_{1},f_{2})\in R^{2}.$
First, we assume $charF\neq 2$. The following two Propositions are from [PPG25].
@@
(v) Let $t>q$ and $\alpha=1$. Then $f\sim(x^{2}+y^{s},y^{t}+exy^{q})$ for a suitable unit $e\in F[[y]]$. If $2t-2q-s\neq 0$ and $p=0$ (or $p\nmid(2t-2q-s)$), then $f\sim(x^{2}+y^{s},y^{t}+xy^{q})$.
@@
Assume $f$ is of the form (3.1). Then if $s\geq 5,t\geq 6$ and $q\geq 3$, then $f$ is of modality at least 2.
@@
Let $g=j_{3}(f)=(x^{2},0).$ Then an open subset in $J_{5}^{\prime}(g)$ is formed by jets equivalent with
$h\sim(x^{2}+y^{5},axy^{3}),$
[…]
For all $a,b\in F$, computation shows $(y^{4},0),(0,y^{4}),(0,y^{5})\notin\widetilde{T}_{h}(\mathcal{K}_{5}h)$. Hence the codimension of $\widetilde{T}_{h}(\mathcal{K}_{5}h)\cap P_{3,5}$ in $P_{3,5}$ is at least 2. By Proposition 2.19, $\mathcal{K}-mod(f)\geq\mathcal{K}_{5}-mod(f)\geq 2$. ∎
@@
By (3.1), $h\sim(x^{2}+\alpha y^{5},y^{6}+axy^{3}+bxy^{4}+cxy^{5})$. One can show, for example, applying $\phi(y)=y-\frac{c}{6}x$, that $h\sim(x^{2}+\alpha y^{5},y^{6}+axy^{3}+bxy^{4})$. This simplifies the computation of the codimension, but does not affect the result.
The following proposition is also from [PPG25].
@@
Let $p=\text{char}(F)$. The following ICIS are the only candidates for being simple (i.e. of modality 0):
@@
    (3.b.ii.2.1) $s\geq 3,\ q\leq 2,$ then $f\sim(xy,x^{2}+y^{m})$ for some $m$ and $q=1$, and $f\sim(x^{2}+y^{m},xy^{2})$ for some $m$ and $q=2$.
@@
$\widetilde{l}_{q,t,t^{\prime}}$ $(x^{2}+y^{4},y^{q+2}+\lambda xy^{q}+uxy^{t}+xy^{t^{\prime}}),\ where$ $\lambda^{2}=-1,q\geq 3,$
@@
     (3.b.i.2.2.3) If p=t=3,s>4,p=t=3,\ s>4, same as the process in [PPG25] Proposition 2.5, $f\sim(x^{2},y^{3}+xy^{2})$, which is simple.
     (3.b.i.2.2.4) If p=t=5,s>4p=t=5,\ s>4, then
$f$ $\sim(x^{2}+y^{s},y^{5}+e(y)xy^{4}),\ where\ e(y)\in F[[x,y]]\ is\ a\ unit$ (3.2)
$\sim(x^{2}+e_{0}y^{s},y^{5}+xy^{4}),\ e_{0}\in F$
@@
where $\widetilde{e}=\frac{1}{1+\frac{e_{0}}{2}y^{s-2}}$. Applying $\phi(x)=\frac{1}{\widetilde{e}}x$ and $\phi(y)=y$ , we have $f\sim(x^{2}+e_{1}y^{2s-1},y^{5}+xy^{4}),$ where $e_{1}=-\frac{\widetilde{e}^{2}e_{0}^{2}}{4}$. Repeating the process, we get $f\sim(x^{2},y^{5}+xy^{4})$ with $p=charF=5$.
@@
   (3.b.ii.1) If $\alpha=0$, by Proposition 3.3, we have $1\leq q\leq 2$ or $q=3,t=5$.
@@
    (3.b.ii.2.2) If $s=3$, then $f\sim(x^{2}+y^{3},xy^{q})$ or $(x^{2}+y^{3},y^{q+2}+xy^{q})$ for $q\geq 3$, which is also simple.
@@
     (3.b.ii.2.3.1) If $t\geq q+4$, then $f\sim(x^{2}+y^{4},y^{t}+e(y)xy^{q}-y^{t-4}(x^{2}+y^{4}))=(x^{2}+y^{4},(e(y)-xy^{t-q-4})xy^{q})\sim(x^{2}+y^{4},xy^{q}),$ where $e(y)=b_{q}+b_{q+1}y+\dots$ is a unit in $F[[x,y]]$ and $b_{q},b_{q+1}$ are defined in (3.1).
     (3.b.ii.2.3.2) If $t=q+3$, then $f\sim(x^{2}+y^{4},y^{q+3}+xy^{q})$ by Proposition 3.2 (v) since in this case 2t2qs=22t-2q-s=2.
     (3.b.ii.2.3.3) If $t=q+2$. Let $g=(x^{2}+y^{4},y^{q+2}+\lambda xy^{q})$, where $\lambda=e(0)\in F^{\times}$, $e(0)$ is the constant term of $e(y)$ in (3.b.ii.2.3.1). Then $g$ is weighted homogeneous of degree 0 with respect to $(a;d)$, where $a=(2,1),d=(4,q+2)$.
@@
for $\lambda^{2}+1=0,t>q$ and $e(y)$ is a unit in $F[[y]]$. In fact, we can show that for some $e(y)$ and $l$, $f\sim(x^{2}+y^{4},y^{q+2}+\lambda xy^{q}+xy^{l})$, while for others, we get a family of unimodal ICIS with the form
$(x^{2}+y^{4},y^{q+2}+\lambda xy^{q}+u(y)xy^{t}+xy^{t^{\prime}}),$ (3.4)
@@
Denote $e_{1}=(x^{2}+y^{4},0),e_{2}=(0,x^{2}+y^{4}),e_{3}=(y^{q+2}+\lambda xy^{q},0),e_{4}=(0,y^{q+2}+\lambda xy^{q}),e_{5}=(2x,\lambda y^{q}),e_{6}=(4y^{3},(q+2)y^{q+1}+\lambda qxy^{q-1})$.
Set the weight $weight(x)=2,weight(y)=1,weight(x^{2}+y^{4})=4,weight(y^{q+2}+\lambda xy^{q})=q+2.$ Then by (2.3), $\widetilde{T}_{g}(F^{1}\mathcal{K}g)$ has the element $x^{i}y^{j}e_{k}$ with:
(a) $k=1,4,\ weight(x^{i}y^{j}e_{k})\geq(5,q+3)$
@@
ProofoftheclaiminProof\ of\ the\ claim\ in\ (3.b.ii.2.3.3.2):
We have $f\sim(x^{2}+y^{4},y^{q+2}+\lambda xy^{q}+e(y)xy^{t})$. If $p\nmid t-q$, we can use an $\alpha,\beta$-trick’ based on the Implicit Function Theorem to show that $f\sim(x^{2}+y^{4},y^{q+2}+\lambda xy^{q}+xy^{t})$ for $t\geq q+1$. See Remark 5.11. For the case $p\mid t-q$, we write
$e(y)=\sum_{i\geq 0}e_{i}y^{q+ip}+\sum_{j\geq 0,\ p\nmid t^{\prime}+j}e^{\prime}_{j}y^{t^{\prime}+j},$
then $f\sim(x^{2}+y^{4},y^{q+2}+\lambda xy^{q}(1+e_{1}y^{p}+e_{2}y^{2p}+\dots)+xy^{t^{\prime}}(e_{0}^{\prime}+e_{1}^{\prime}y+e_{2}^{\prime}y^{2}+\dots))$ with $p\nmid t^{\prime}-q$. Therefore we can use $\alpha,\beta$-trick again to reduce $e_{0}^{\prime}+e_{1}^{\prime}y+\dots$. Then we have $f\sim(x^{2}+y^{4},y^{q+2}+\lambda xy^{q}+u(y)xy^{t}+xy^{t^{\prime}})$, where $u(y)=u_{0}+u_{1}y^{p}+u_{2}y^{2p}+\dots$ as we want.